OpenStudy (anonymous):

\[\lim_{(x,y) \to (0,0)} x\ cos\left ({1 \over y} \right )\]

OpenStudy (jamesj):

Question for you. What can you say about the size of cos(1/y)?

OpenStudy (anonymous):

it goes to 1 as y goes to 0?

OpenStudy (jamesj):

In general, what can you say about the function cos(theta), what is its max and min?

OpenStudy (anonymous):

absolute value <= 1

OpenStudy (jamesj):

yes. Hence what can be bound \[ \left| x \cos(1/y) \right| \] by?

OpenStudy (anonymous):

|x cos(1/y)| <= |x|

OpenStudy (jamesj):

Yes. Hence the limit is equal to what?

OpenStudy (anonymous):

by squeeze theorem, the limit is 0?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

Awesome, thank you